OpenStudy (anonymous):

Solve for the variable: -2v – 3 = 25 A. -14 B. 7 C. 1 D. 11

OpenStudy (anonymous):

A) -14 -2v-3=25 -2v=28 -v=14 v=-14

OpenStudy (anonymous):

A. -14 because a negative times a negative must be a positive

OpenStudy (anonymous):

im very grateful for your guy's help thank you

OpenStudy (anonymous):

your welcome... Need anything else

OpenStudy (anonymous):

ill make sure to post any future questions :)

OpenStudy (anonymous):

ok just ask me and ill be happy to help :)